OpenStudy (anonymous):

When fully loaded, a 747 airplane flies at 475.13 miles per hour. How far can a loaded 747 travel in 5.25 hours? Round your answer to the nearest hundredth.

OpenStudy (anonymous):

25000

OpenStudy (anonymous):

2500

OpenStudy (texaschic101):

475.13/1 = x/5.25 (474.13 miles per hr to x miles per 5.25 hrs cross multiply x = (475.13)(5.25) x = 2494.43 rounded = 2500

OpenStudy (anonymous):

can any one help me with this Gino designed a garden for his rectangular backyard. The design he drew on paper is 7 inches wide. The width of his backyard is forty-two feet. If his drawing is to scale and is 8 inches long, how many feet long is his backyard? Write a proportion using words. Then, use the proportion to calculate the yard length. Show your work.

OpenStudy (anonymous):

@Cashmoney8633533 Post your own question.